2) Suppose you are presented with the following table, where all figures are rendered in billions of dollars: MMMF Reserves S all of them please <3
Answer #1

(a)

Calculate monetary base in 2017 -

Monetary base = Currency + Reserves

Monetary base = 1,270 + 2,400

Monetary base = 3,570

The monetary base in 2017 is 3,570.

Calculate monetary base in 2018 -

Monetary base = Currency + Reserves

Monetary base = 1,330 + 2,500

Monetary base = 3,830

The monetary base in 2018 is 3,830.

(b)

Calculate M1 in 2017 -

M1 = Currency + Demand Deposit

M1 = 1,270 + 1,200

M1 = 2,470

The M1 in 2017 is 2,470

Calculate M1 in 2018 -

M1 = Currency + Demand Deposit

M1 = 1,330 + 1,220

M1 = 2,550

The M1 in 2018 is 2,550

(c)

Calculate M2 in 2017 -

M2 = Currency + Demand Deposit + Saving Deposit + Small Time Deposit + MMMF

M2 = 1,270 + 1,200 + 7,700 + 505 + 600

M2 = 11,275

The M2 in 2017 is 11,275

Calculate M2 in 2018 -

M2 = Currency + Demand Deposit + Saving Deposit + Small Time Deposit + MMMF

M2 = 1,330 + 1,220 + 8,200 + 602 + 720

M2 = 12,072

The M2 in 2018 is 12,072

(d)

M1 in 2017 = 2,470

M2 in 2018 = 2,550

Calculate the growth rate of M1 -

Growth rate = [(M1 in 2018 - M1 in 2017)/M1 in 2017] * 100

Growth rate = [(2,550 - 2,470)/2,470] * 100

Growth rate = 3.24%

The growth rate of M1 is 3.24%

M1 is basis for monetary growth rate. This means that the growth rate of M1 is equal to growth rate of money supply.

So, growth rate of money supply is 3.24%.

Velocity is constant.

Growth rate of real GDP is 1%.

Calculate the rate of inflation -

Rate of inflation = Growth rate of money supply - Growth rate of real GDP

Rate of inflation = 3.24% - 1% = 2.24%

The rate of inflation is 2.24%

(e)

Annual nominal interest rate = 5%

Inflation rate = 2.24%

Real interest rate = Annual nominal interest rate - Inflation rate = 5% - 2.24% = 2.76%

Calculate the real interest return on deposit -

Real interest return = Amount of deposit * Real interest rate

Real interest return = $1,000 * 0.0276 = $27.6

So,

The real interest return on deposit will be $27.6